OpenStudy (anonymous):

If 95 is added to a number, the result is 37 less than three times the number. Find the number.

OpenStudy (anonymous):

x+95 = 3x-37 132 = 2x x = 66

OpenStudy (anonymous):

n+95=3n-37 95=2n-37 2n=132 n=66

OpenStudy (anonymous):

Let x be the number So x + 95 = 3x-37 add 37 to both sides x + 132= 3x subtract x from both sides 132= 2x divied both by 2 x = 66 Voila :)